copper wire of length 150 cm is heated from 30°C to 130°C. If its length gets increased by 2.55 mm,
the coefficient of linear expansion of copper is

Grade:12th Pass

2 Answers

Sudheesh Singanamalla
13 years ago

Length of copper wire = 150 cm = 1500 mm

Temperature initial (Ti) = 30

Temperature Final (Tf) = 130

change in temperature = 130 - 30 = 100

change in  length = 2.55 mm

change in length / actual length = coefficient of linear expannsion * Change in temperature.

2.55/(1500 * 100 ) = coefficient of linear expansion.

coefficient of linear expansion. = 0.000017 = 1.7 * 10 ^ -5

----------------------------------------------------------------------------------------

Please approve !!

---------------------------------------------------------------------------------------

Ritvik Gautam
13 years ago

We know that the change in length of wire is,

l'=lαT

where l' is the change in length

l is original length

α is the coefficient of linear expansion

and T is the change in temperature.

Substituting the values,

2.55x10-3 = 0.150 x α x 100

Hence, we get

α=1.5x10-4

 

 

Please Approve!!